OpenStudy (anonymous):

The distance, d, in kilometres, travelled by a plane after t hours can be represented by d(t)=-4t^3=40t^2+500t, Where t is greater than or equal to 0 and less than or equal to 10. How long does the plane take to fly 4088km?

OpenStudy (phi):

is there a typo in your equation?

OpenStudy (anonymous):

yes, d(t)=-4t^3+40t^2+500t

OpenStudy (phi):

they want you to solve for "t" when \[ -4t^3+40t^2+500t=4088 \] or \[ -4t^3+40t^2+500t-4088=0 \]

OpenStudy (phi):

we can simplify that equation a little by dividing both sides by -4 \[ t^3 -10t^2 -125t +1022 = 0 \]

OpenStudy (anonymous):

and then what

OpenStudy (phi):

we can graph it

OpenStudy (phi):

What are you studying?

OpenStudy (anonymous):

Advanced functions and solving polynomials

OpenStudy (phi):

This is a cubic, and though cubics have a "closed formula", it is a bit hairy. Are you supposed to use that formula? i.e. how do they expect you to solve this?

OpenStudy (anonymous):

i have no idea ! lol

OpenStudy (phi):

Are you saying you never saw an example of solving a cubic?

OpenStudy (anonymous):

we are taught to factor it and find the zeros